# Motion Before Action: Diffusing Object Motion as Manipulation Condition

### v1

@@ Abstract
Inferring object motion representations from observations enhances the performance of robotic manipulation tasks. This paper introduces a new paradigm for robot imitation learning that generates action sequences by reasoning about object motion from visual observations. We propose MBA, a novel module that employs two cascaded diffusion processes for object motion generation and robot action generation under object motion guidance. MBA first predicts the future pose sequence of the object based on observations, then uses this sequence as a condition to guide robot action generation. Designed as a plug-and-play component, MBA can be flexibly integrated into existing robotic manipulation policies with diffusion action heads. Extensive experiments in both simulated and real-world environments demonstrate that our approach substantially improves the performance of existing policies across a wide range of manipulation tasks.
Index Terms:
Robotic Manipulation, Imitation Learning
@@ I Introduction
In contrast to the biological approach, most existing robot policies [ACT, octo, DP, RISE, cage, DP3] are predominantly guided by observation, employing feature encoder and adopting generative approaches to predict actions. While effective, it often results in an overreliance on environmental cues, with the model focusing on memorizing observation features rather than reasoning about object motion patterns, as humans do. Consequently, when encountering extensive pose shifts in real-world objects or actions, many policies often struggle to generalize effectively [octo, openvla], which can limit their practical performance.
@@
In this work, we introduce Motion Before Action (referred to as MBA), a novel module that equips existing robotic manipulation policies with such reasoning skills. We observe that representing the object motion using 6D pose aligns with the robot’s end-effector pose in the same format, making the actions compatible with this representation. Consequently, they exhibit mathematical consistency. During task execution, at each time step, the robot pose, object pose, and action are in proximity within the same space, with their kinematic relationships following a learnable pattern, demonstrating physical consistency [xu2021end]. Based on these two distributional consistencies, we hypothesize the learnability of object pose representation in diffusion models [DDPM, DDIM]: 1) diffusion models have been shown to effectively reconstruct actions [DP], so they should similarly excel at reconstructing analogous object poses; 2) just as the robot’s pose can serve as a condition [octo] for generating actions via diffusion models, so does the object pose.
@@
  • We introduce MBA, a flexible auxiliary module that can be easily integrated into existing policies with diffusion heads, rather than functioning as an independent policy.
  • We conduct comparative experiments with MBA on three 2D and 3D robotic manipulation policies with diffusion action heads [DP, DP3, RISE], demonstrating substantial performance improvements across various tasks. These tasks, comprising 57 tasks from 3 simulation benchmarks and 3 real-world tasks, involve articulated object manipulation, soft and rigid body manipulation, tool use, non-tool use, and diverse action patterns. Results show that MBA consistently enhances the performance of such policies in both simulated and real-world environments.
@@ II-B Object-Centric Manipulation Learning
To overcome these limitations, inspired by the success of flow-based methods [teed2020raft, goyal2022ifor], recent work has utilized object flow as a general affordance for guiding manipulation policies [WenAnyPoint, VecerikRobotap, XuFlowCrossDomain, BharadhwajTrack2act, YuanGeneralFlow, EisnerFlowbot3d, ZhangFlowbotPlusPlus, SeitaToolflownet]. These policies map vision to motion by anchoring visual features and constructing object flow, which is then translated into robot actions using either heuristic transformations [YuanGeneralFlow, BharadhwajTrack2act, ZhangFlowbotPlusPlus, SeitaToolflownet, EisnerFlowbot3d] or learned models [WenAnyPoint, XuFlowCrossDomain]. Many methods incorporate object pose as an intermediate representation [YuanGeneralFlow, BharadhwajTrack2act, SeitaToolflownet] in such flow-to-action translations. However, using flow as an indirect motion representation can introduce ambiguity due to the vision-motion gap [WenAnyPoint, XuFlowCrossDomain, BharadhwajTrack2act], complicating flow-to-action learning. Moreover, defining such heuristic transformations is challenging, often hindering generalization in dynamic or novel environments.
@@ III Method
Figure 1: Overview of MBA. MBA takes the current observation as input, which could be in the form of 3D point clouds or RGB images from different viewpoints. Object pose sequences are sampled as actions with denoising diffusion from the object policy to be part of the framework’s output. Conditioning on the observations and object pose actions, MBA samples deployable robot actions with denoising diffusion from the robot policy. These actions are executed within the workspace to update the environment state and the observations.
MBA is an object motion generation module $p(\mathbf{M}|\mathbf{O})$ that is readily compatible with any existing policy with diffusion action heads, enhancing versatility and efficiency across various robotic manipulation contexts. Given an observation, the perception module of the original policy first processes it into observation features. These observation features are then passed to MBA, where they serve as a condition for the object pose sequence denoising process. Subsequently, the diffusion action head of the original policy takes both the observation features and the denoised object pose sequence as inputs and outputs the corresponding robot actions. During training, both ground-truth object pose sequences and action sequences are used as supervision for the predicted object pose sequences from MBA and action sequences from the action head. At inference time, the policy with MBA operates in an end-to-end manner, without requiring ground-truth object poses. The full execution framework of MBA is illustrated in Fig. 1.
@@ III-A Object Motion Generation
As discussed in §II-B, we opt for object pose as a direct representation to describe object motion. The object pose $\mathrm{m}_{t}$ is represented as a 9D vector, combining 3D translation and 6D rotation [rot6d], to align with robot action representations [DP, DP3, RISE]. At each observation step $t$, we use the observation features $O_{t}$ as the condition to the diffusion process [DDPM], which denoises the object pose sequence for the next $T_{m}$ steps $\mathbf{M}_{t:t+T_{m}}=(\mathrm{m}_{t},\mathrm{m}_{t+1},\mathrm{m}_{t+2}...\mathrm{m}_{t+T_{m}})$.
@@ III-B Robot Action Generation under Object Motion Guidance
Following [DP, DP3, RISE], the robot action $\mathrm{a}_{t}$ is defined by the end-effector pose and gripper width, using a 10D representation that concatenates 3D translation, 6D rotation [rot6d], and 1D gripper width together. Similarly, we use the diffusion process to generate the action sequence at the next $T_{a}$ steps: $\mathbf{A}_{t:t+T_{a}}=(\mathrm{a}_{t},\mathrm{a}_{t+1},\mathrm{a}_{t+2}...\mathrm{a}_{t+T_{a}})$.
@@
III-C Execution
Algorithm 1 MBA Execution
@@ IV Simulation Experiments
TABLE I: Average success rate and standard deviation of policies integrating MBA compared with baselines under 3 random seeds for 57 simulation tasks.
@@ IV-A Setup
Baselines. The focus of this work is to demonstrate that the introduction of MBA as a module can universally enhance the performance of existing policies with diffusion heads. Therefore, in our simulation experiments, we select representative Diffusion Policy [DP] and 3D Diffusion Policy [DP3] as our 2D and 3D baseline respectively. We then integrate the MBA module into these baselines and compare the performance. To ensure fairness in the experiments, we ensure that MBA and the corresponding baseline methods use the same expert demonstrations for training, with an equal amount of training steps. During the execution phase, both methods undergo the same number of observation and inference steps.
@@
Protocols. Following [DP3], We conduct 3 runs for each experiment, using seed numbers 0, 1, and 2. For each seed, we evaluate 20 episodes every 200 training epochs and compute the average of the top 5 success rates. We then report the mean and standard deviation of success rates across the 3 seeds.
IV-B Results
@@
TABLE II: Detailed performance of policies integrating MBA and baselines in several typical simulation tasks.
@@
Figure 2: The relationship between the success rates of MBA and baselines and the number of training steps in a single experiment.
Integrating MBA results in more stable and superior performance (Q1). We report the average success rates and standard deviations (across three random seeds) for all simulation tasks in the Adroit, Dexart, and MetaWorld environments, across four difficulty levels, in Tab. I. On average, MBA outperforms the baseline in the majority of benchmarks, achieving a 14.2% increase in the average success rate over DP and a 6.2% increase over DP3. Additionally, we observe a general reduction in the average standard deviation of task execution across all benchmarks, further confirming the robustness of MBA. The results also demonstrate that MBA significantly improves success rates in tasks with higher difficulty levels. Detailed examples are provided in Tab. II, where we report the average success rate and standard deviation for particular tasks. Notably, in tasks that require precise contact and fine manipulation within narrow action spaces, the accurate object pose estimation and motion prediction powered by MBA can substantially enhance the execution capability and robustness of the robotic manipulation policies.
@@ V-A Setup
Platform. We employ a Flexiv Rizon robotic arm with a Robotiq 2F-85 gripper to manipulate objects. A global Intel RealSense D415 RGB-D camera is positioned in front of the robot for 3D perception, capturing single-view workspace point clouds. The robot workspace is defined as a 40 cm $\times$ 60 cm rectangular area in front of the robot. The overview of the real-world robot platform is shown in Fig. 3. All devices are connected to a workstation with an Intel i9-10980XE CPU and an NVIDIA 2080 Ti GPU for data collection and evaluation. Since the training process of MBA requires 6D object motion data for supervision, we established a Motion Capture (MoCap) system comprising five OptiTrack Prime 13W infrared cameras. Before the training phase, we inpaint those MoCap markers within RGB-D observations to maintain visual consistency between the training and deployment. Notably, no motion capture is needed during deployment.
Tasks. We select three tasks for experimentation: Cut Clay (a tool-use contact-rich task), and Put Bread into Pot (soft body manipulation), and Open Drawer (articulated object manipulation), as illustrated in Fig. 3.
@@
Baseline. We select RISE [RISE] as the baseline, which is a state-of-the-art real-world robotic manipulation policy that relies solely on 3D scene input and generates actions through a diffusion action head, conditioned on 3D features extracted via sparse convolutional networks [spatioenc] and transformers [transformer]. We then integrate the MBA module to predict object pose sequences prior to the action head of RISE, with additional conditioning on the generated object motion.
Protocols. In the real-world evaluation, we conduct 20 trials per method for each task, unless stated otherwise. All methods are compared under nearly identical randomized initial scene configurations for each trial.
@@ V-B Cut Clay
In this experiment, we randomize the position and orientation of the cutting board, knife holder, and foam mat for each test. Special attention is given to varying the relative positions of these three objects to thoroughly assess policy robustness. Additionally, the shape of the clay is changed each time to meet generalization requirements for real-world applications. We define five evaluation metrics for this task:
@@
Methods Metrics
[…]
RISE $95$ $70$ $30$ $50$ $11$
RISE w. MBA $100$ $90$ $55$ $65$ $12$
TABLE III: Experimental results of the Cut Clay task.
From Tab. III, we observe that RISE with MBA shows significant improvements over RISE in all stages of task execution, particularly in the cutting and separation stage. This improvement is attributed to the need for rotating the cutting surface and splitting it when the knife cuts into the clay, which is inherently a 6-DoF task. MBA leverages the prediction of the tool motion to provide feedback on the execution of the current action, resulting in superior task performance compared to the baseline. However, MBA still exhibits repeated attempts to grasp the knife. We hypothesize that this issue arises from estimation errors in this task, where the thinness of the blade amplifies these errors, preventing them from being negligible.
Figure 3: Overview of our real-world robot platform and the illustration of three real-world tasks.
@@ V-C Put Bread into Pot
The Put Bread into Pot task is a classic pick-and-place task, where the objective is to pick the bread from a bowl and place it into a pot. We chose this task because bread is a soft object. Unlike rigid bodies, soft objects can deform significantly due to the gripper’s hold and the pressure from touching the bottom of the pot, which affects the object’s pose, increasing the challenge for MBA in predicting the object pose sequence. We aim to assess the robustness and generalization ability of MBA through this task.
@@
Method Metrics
Success Rate (%) $\uparrow$ Attempt $\downarrow$
RISE $80$ $7$
RISE w. MBA $95$ $1$
TABLE IV: Experimental results of the Put Bread into Pot task.
The results, as shown in Tab. IV, indicate that RISE with MBA outperforms vanilla RISE by 15% in average success rate. In most cases, RISE with MBA successfully grasps the bread on the first attempt, whereas RISE often requires multiple attempts. This demonstrates that MBA can effectively handle object pose sequence prediction for soft objects, contributing to more accurate object localization and grasping.

### v3

@@ Abstract
Inferring object motion representations from observations enhances the performance of robotic manipulation tasks. This paper introduces a new paradigm for robot imitation learning that generates action sequences by reasoning about object motion from visual observations. We propose MBA, a novel module that employs two cascaded diffusion processes for object motion generation and robot action generation under object motion guidance. MBA first predicts the future pose sequence of the object based on observations, and then uses this sequence as a condition to guide robot action generation. Designed as a plug-and-play component, MBA can be flexibly integrated into existing robotic manipulation policies with diffusion action heads. Extensive experiments in both simulated and real-world environments demonstrate that our approach substantially improves the performance of existing policies across a wide range of manipulation tasks. Project page: https://selen-suyue.github.io/MBApage/
@@ I Introduction
In contrast to the biological approach, most existing robot policies [ACT, octo, DP, RISE, cage, DP3] are predominantly guided by observation, employing feature encoders and adopting generative approaches to predict actions. While effective, it often results in an overreliance on environmental cues, with the model focusing on memorizing observation features rather than reasoning about object motion patterns, as humans do. Consequently, when encountering extensive pose shifts in real-world objects or actions, many policies often struggle to generalize effectively [octo, openvla], which can limit their practical performance.
@@
Figure 1: Understanding object motion before action leads to better manipulation. Unlike existing methods that predict actions directly from observations, our approach first infers future object motions, enabling more accurate and goal-driven action prediction.
In this work, we introduce Motion Before Action (referred to as MBA), a novel module that equips existing robotic manipulation policies with such reasoning skills. We observe that representing the object motion using 6D pose aligns with the robot’s end-effector pose in the same format, making the actions compatible with this representation. Consequently, they exhibit mathematical consistency. During task execution, at each time step, the robot pose, object pose, and action are in proximity within the same space, with their kinematic relationships following a learnable pattern, demonstrating physical consistency [xu2021end]. These consistencies suggest that the object poses and the robot poses and actions share similar distributions. The ability of probabilistic models, exemplified by diffusion models, to learn such similar distributions is transferable, as indicated in [DDPM, DDIM]. Therefore, we propose that object poses can be generated by the diffusion process in the same way as actions. These object poses can also effectively guide the action denoising process as conditions, just as robot poses do.
Building on this foundation, we design MBA as a cascade of two diffusion modules. MBA seamlessly integrates with any existing diffusion-based policy. The observation features encoded by the given policy serve as conditional inputs to the first diffusion module, which predicts future object pose sequences. These predicted pose sequences, combined with the initial observations, are then fed as joint conditions to the diffusion action head, effectively guiding the policy’s action generation.
We conduct comparative experiments with MBA on three 2D and 3D robotic manipulation policies with diffusion action heads [DP, DP3, RISE], demonstrating substantial performance improvements across various tasks. These tasks, comprising 57 tasks from 3 simulation benchmarks and 4 real-world tasks, involve articulated object manipulation, soft and rigid body manipulation, tool use, non-tool use, and diverse action patterns. Results show that MBA consistently enhances the performance of such policies in both simulated and real-world environments.
@@
  • We introduce MBA, a flexible auxiliary module that can be easily integrated into existing policies with diffusion heads, rather than functioning as an independent policy. Extensive experiments demonstrate its ability to significantly improve the performance of these policies, highlighting its broad applicability and potential to enhance a wide range of robotic manipulation tasks.
@@ II-B Object-Centric Manipulation Learning
To overcome these limitations, inspired by the success of flow-based methods [teed2020raft, goyal2022ifor], recent work has utilized object flow as a general affordance for guiding manipulation policies [ATM, VecerikRobotap, XuFlowCrossDomain, BharadhwajTrack2act, YuanGeneralFlow, EisnerFlowbot3d, ZhangFlowbotPlusPlus, SeitaToolflownet]. These policies map vision to motion by anchoring visual features and constructing object flow, which is then translated into robot actions using either heuristic transformations [YuanGeneralFlow, BharadhwajTrack2act, ZhangFlowbotPlusPlus, SeitaToolflownet, EisnerFlowbot3d] or learned models [ATM, XuFlowCrossDomain]. Many methods incorporate object pose as an intermediate representation [YuanGeneralFlow, BharadhwajTrack2act, SeitaToolflownet] in such flow-to-action translations. However, using flow as an indirect motion representation can introduce ambiguity due to the vision-motion gap [ATM, XuFlowCrossDomain, BharadhwajTrack2act], complicating flow-to-action learning. Moreover, defining such heuristic transformations is challenging, often hindering generalization in dynamic or novel environments.
@@ III Method
Figure 2: Overview of MBA pipeline. MBA takes the current observation as input, which could be in the form of 3D point clouds or RGB images from different viewpoints. Object pose sequences are sampled as actions with denoising diffusion from the object policy to be part of the framework’s output. Conditioning on the observations and object pose actions, MBA samples deployable robot actions with denoising diffusion from the robot policy. These actions are executed within the workspace to update the environment state and the observations.
MBA has an object motion generation module $p(\mathbf{M}|\mathbf{O})$ that is readily compatible with any existing policy with diffusion action heads, enhancing versatility and efficiency across various robotic manipulation contexts. Given an observation, the perception module of the original policy first processes it into observation features. These observation features are then passed to MBA, where they serve as a condition for the object pose sequence denoising process. Subsequently, the diffusion action head of the original policy takes both the observation features and the denoised object pose sequence as inputs and outputs the corresponding robot actions. During training, both ground-truth object pose sequences and action sequences are used as supervision for the predicted object pose sequences from MBA and action sequences from the action head. At inference time, the policy with MBA operates in an end-to-end manner, without requiring ground-truth object poses. The full execution framework of MBA is illustrated in Fig. 2.
@@ III-A Object Motion Generation
As discussed in §II-B, we opt for object pose as a direct representation to describe object motion. The object pose $\mathrm{m}_{t}$ is represented as a 9D vector, combining 3D translation and 6D rotation [rot6d], to align with robot action representations [DP, DP3, RISE]. At each observation step $t$, we use the observation features $O_{t}$ as the condition to the diffusion process [DDPM], which denoises the object pose sequence for the next $T_{m}$ steps $\mathbf{M}_{t:t+T_{m}}=(\mathrm{m}_{t},\mathrm{m}_{t+1},\mathrm{m}_{t+2}...\mathrm{m}_{t+T_{m}-1})$.
@@ III-B Robot Action Generation under Object Motion Guidance
Following [DP, DP3, RISE], the robot action $\mathrm{a}_{t}$ is defined by the end-effector pose and gripper width, using a 10D representation that concatenates 3D translation, 6D rotation [rot6d], and 1D gripper width together. Similarly, we use the diffusion process to generate the action sequence at the next $T_{a}$ steps: $\mathbf{A}_{t:t+T_{a}}=(\mathrm{a}_{t},\mathrm{a}_{t+1},\mathrm{a}_{t+2}...\mathrm{a}_{t+T_{a}-1})$.
@@
Algorithm 1 MBA Execution Process
@@ IV Simulation Experiments
TABLE I: Performance comparison of policy models with and without MBA integration across 57 simulation tasks. Average success rates ($\pm$ standard deviation) are reported under three random seeds.
@@ IV-A Setup
Baselines. The focus of this work is to demonstrate that the introduction of MBA as a module can universally enhance the performance of existing policies with diffusion heads. Therefore, in our simulation experiments, we select representative Diffusion Policy (DP[DP], 3D Diffusion Policy (DP3[DP3] as our 2D and 3D baselines. We then integrate the MBA module into these baselines and compare the performance. To ensure fairness in the experiments, we ensure that MBA and the corresponding baseline methods use the same expert demonstrations for training, with an equal amount of training steps. During the execution phase, both methods undergo the same number of observation and inference steps.
@@
Protocols. Following [DP3], We conduct 3 runs for each experiment, using seed numbers 0, 1, and 2. For each seed, we evaluate 20 episodes every 200 training epochs as a test node, compute the average success rate over 20 episodes for each test node, and then average the top 5 performing test nodes. This accounts for the inherent instability of imitation learning policies during training in simulated environments, where different policies converge at varying training stages. We then report the mean and standard deviation of success rates across the 3 seeds.
@@
TABLE II: Task-level success rates across 11 simulation tasks in 3 environments comparing MBA-augmented and baseline policies.
@@
Figure 3: Average learning curves (success rate - training steps) over three runs comparing MBA-augmented and baseline policies.
IV-B Results
Integrating MBA results in more stable and superior performance (Q1). We report the average success rates and standard deviations (across three random seeds) for all simulation tasks in the Adroit, Dexart, and MetaWorld environments, across four difficulty levels, in Table. I. On average, MBA outperforms the baseline in the majority of benchmarks, achieving a 14.2% increase in the average success rate over DP and a 6.2% increase over DP3. Additionally, we observe a general reduction in the average standard deviation of task execution across all benchmarks, further confirming the robustness of MBA. The results also demonstrate that MBA significantly improves success rates in tasks with higher difficulty levels. Detailed examples are provided in Table. II, where we report the average success rate and standard deviation for particular tasks. Notably, in tasks that require precise contact and fine manipulation within narrow action spaces, MBA can enhance the execution capability and robustness of the robotic manipulation policies. We attribute this success primarily to two factors. First, when objects are stationary, MBA’s accurate prediction (can also be regarded as a pose estimation in this situation) of object pose enables the robot to localize and grip the objects effectively. Second, when objects are in motion, the strong correlation between the predicted object pose sequences and the robot’s action sequences provides robust guidance and calibration for action generation.
@@ V-A Setup
Platform. We employ a Flexiv Rizon robotic arm with a Robotiq 2F-85 gripper to manipulate objects. A global Intel RealSense D415 RGB-D camera is positioned in front of the robot for 3D perception, capturing single-view workspace point clouds. The robot workspace is defined as a 40 cm $\times$ 60 cm rectangular area in front of the robot. The overview of the real-world robot platform is shown in Fig. 4. All devices are connected to a workstation with an Intel i9-10980XE CPU and an NVIDIA 2080 Ti GPU for data collection and evaluation. Since the training process of MBA requires 6D object motion data for supervision, we set up a Motion Capture (MoCap) system comprising five OptiTrack Prime 13W infrared cameras. Notably, no motion capture system or markers are required during deployment.
Figure 4: Real-world deployment platform and execution process of four manipulation tasks.
Tasks. We select 4 tasks for experimentation: Cut Clay (a tool-use contact-rich task), Put Bread into Pot (soft body manipulation), Open Drawer (articulated object manipulation), and Pour Balls (6-DoF task) as illustrated in Fig. 4.
@@
Baseline. We retain DP and DP3 from the previous section as baselines, and additionally include RISE [RISE], a SOTA real-world policy model that operates solely on 3D scene input. RISE generates actions through a diffusion head, conditioned on 3D features extracted via sparse convolution networks [spatioenc] and transformers. We integrate the MBA module into these three policies, placing it before the action head to predict the future object pose sequence as the action condition.
Protocols. In the real-world evaluation, we conduct 20 trials per method for each task unless stated otherwise. All methods are compared under nearly identical randomized initial scene configurations for each trial.
@@ V-B Cut Clay
In this experiment, we randomize the position and orientation of the cutting board, knife holder, and foam mat for each test. Special attention is given to varying the relative positions of these three objects to assess policy robustness thoroughly. Additionally, the shape of the clay is changed each time to meet the generalization requirements for real-world applications. We define five evaluation metrics for this task:
@@
Method Cut Clay Put Bread into Pot Open Drawer Pour Balls
Pick (%) Cut (%) Sep. (%) Place (%) Attempt $\downarrow$ Succ. (%) $\uparrow$ Attempt $\downarrow$ Succ. (%) $\uparrow$ Pour (%) Balls $\uparrow$ Pick (%)
RISE [RISE] 95 70 30 50 11 80 7 37.5 85 8.15 85
RISE w. MBA 100 90 55 65 12 95 1 52.5 100 9.60 100
DP3 [DP3] 95 35 10 15 53 25 16 20 55 1.20 60
DP3 w. MBA 80 60 20 25 25 45 29 55 55 1.65 60
DP [DP] - 40 12 20 10 0.80 30
DP w. MBA - 45 16 30 40 3.60 60
TABLE III: Performance comparison of policy models with and without MBA integration across four real-world tasks.
We observe limitations with 2D policies on this long-horizon multi-object task as the appearance of clay is similar in 2D visual observations before and after the cutting process. Therefore, we restrict our comparison to 3D policies. From Table. III, we observe that RISE with MBA shows significant improvements over RISE in all stages of task execution, particularly in the cutting and separation stage. A similar improvement is observed with DP3. This improvement is attributed to the need for rotating the cutting surface and splitting it when the knife cuts into the clay, which is inherently a 6-DoF task. MBA leverages the prediction of the tool motion to provide feedback on the execution of the current action, resulting in superior task performance compared to the baseline. However, MBA still exhibits repeated attempts to grasp the knife. We hypothesize that this issue arises from estimation errors in this task, where the thinness of the blade amplifies these errors.
@@ V-C Put Bread into Pot
The Put Bread into Pot task is a classic pick-and-place task, where the objective is to pick the bread from a bowl and place it into a pot. We choose this task because bread is a soft object. Unlike rigid bodies, soft objects can deform significantly due to the gripper’s hold and the pressure from touching the bottom of the pot, which affects the object’s pose, increasing the challenge for MBA in predicting the object pose sequence. We aim to assess the robustness and generalization ability of MBA through this task.
@@
The results, as shown in Table. III, indicate that RISE with MBA outperforms vanilla RISE by 15% in average success rate. In most cases, RISE with MBA successfully grasps the bread on the first attempt, whereas RISE often requires multiple attempts. This demonstrates that MBA can effectively handle object pose sequence prediction for soft objects, contributing to more accurate object localization and grasping.
Integrating MBA with both DP and DP3 also notably improves success rates. However, we also observe that a higher error rate in the base policy — primarily caused by the visual encoder — leads to increased attempts by MBA. This is likely due to the error accumulates when the visual backbone has a significant error, necessitating more attempts to compensate.